OpenStudy (anonymous):

Mersenne primes. So is this true or false, that if n is prime then (2^n -1) must also be prime?

OpenStudy (freckles):

hmmm I don't think so (could be wrong) maybe we can try to find a counterexample

OpenStudy (freckles):

n=2 we have M=3 n=3 we have M=7 n=5 we have M=31 n=7 we have M=127 so far these M's have been prime try the next prime number n

OpenStudy (freckles):

do you think 2^(11) -1 is prime or not?

OpenStudy (anonymous):

yes, maybe...............

OpenStudy (anonymous):

11 is prime

OpenStudy (anonymous):

2047 does not divide 2,3,5,7,11,13,17, or 19.

OpenStudy (freckles):

2047 is composite

OpenStudy (freckles):

and I think you mean the other way around

OpenStudy (freckles):

but yeah 23 and 89 both divide 2047

OpenStudy (anonymous):

n = 13; 8191 is prime n = 17; 131071 is prime n = 19; 524287 is prime Something is wrong........... 2^11-1 should be prime.........

OpenStudy (freckles):

why is something wrong?

OpenStudy (freckles):

you found a counterexample which means the statement above is false

OpenStudy (freckles):

n being prime doesn't imply M_n is prime

OpenStudy (anonymous):

is 8388607 = 2^23-1 prime

OpenStudy (freckles):

That is way bigger than 2047 lol I would have to use a computer tell me

OpenStudy (anonymous):

i divide by numbers 1-30 none of them work. if this number is composite then i believe this statement is false

OpenStudy (freckles):

we already found a counterexample so you should already believe the statement is false

OpenStudy (freckles):

is 11 not prime? is 2^(11)-1 not 2047 which is the product of 23 and 89?

OpenStudy (anonymous):

ok, I am just wondering now if that is the only correction to this statement......

OpenStudy (freckles):

nope

OpenStudy (freckles):

there are probably infinitely more

OpenStudy (freckles):

http://en.wikipedia.org/wiki/Mersenne_prime

OpenStudy (freckles):

they give more counterexamples there

OpenStudy (anonymous):

i got the link

OpenStudy (freckles):

like notice in there n list they skip 11,23,29,...

OpenStudy (freckles):

37,41, and so on

OpenStudy (freckles):

http://www.wolframalpha.com/input/?i=factors+of+%282%5E%2823%29-1%29 for your question earlier

OpenStudy (freckles):

as you can see for n=23 M was composite since 2^(23)-1 =47(178481)

OpenStudy (freckles):

no wonder you didn't find the divisors they were pretty big

OpenStudy (anonymous):

ok, biggest prime ever is 47983739879261

OpenStudy (freckles):

there should be bigger primes than that right?

OpenStudy (anonymous):

2^.....................-1

OpenStudy (anonymous):

(893041296^28.e+1380219695) + (3820981901^697910283685247e+3789281 +43) + 3810947265444810394765539821317422670881596

OpenStudy (anonymous):

is prime